# HARNESSING Nth ROOT GATES FOR ENERGY STORAGE

### v1

@@ VIII Conclusion
E.F. acknowledges support from EPSRC
@@
Figure 10: Ergotropy $W_{max}$ of the total system for $0\ \leq$ Iterations $\ \leq\ 2N$. For each panel $N$ is indicated on the x-axis; protocol corresponding to case 3 with three-qubit systems. Note each panel corresponds to 2 cycles of iterations. Parameters: $\epsilon_{1j}=0,\ \epsilon_{2j}=1$ for $j=A,B$, and $C$.
Fig. 10 shows how changing $N$ affects the ergotropy for case 3 and the three-qubit system. We observe a ”smoothing” of the evolution of the ergotropy for increasing $N$.
@@
where $T_{\alpha}=\alpha/M$ is the process duration up to the $\alpha$ iteration and $M$ is the total number of iterations considered (here $M=30$ corresponding to two cycles). For $\alpha=M$, $T_{\alpha}=1$. $\langle W_{\alpha}\rangle$ is the average work defined as $\langle W_{\alpha}\rangle=E^{\alpha}-E^{i}$, where $E^{\alpha}=Tr\left[\rho^{\alpha}H\right]$.
When evaluating the power of a quantum battery there is more depth than just the difference in average internal energy of the system. While the average internal energy plays a role, this is not a one to one comparison with the ergotropy, which as thoroughly discussed in the main article, is a better measure for quantum battery performance. With that in mind we introduce,
@@
where $P_{\Delta W_{max}}$ is the power of charging or discharging and $\Delta W_{max}$ is the variation in ergotropy up to the $\alpha$ iteration. Equation 13 characterises the rate at which the system generates ergotropy. The total ergotropy is not used as we are interested in charging from some initial state ($\rho^{i}$) that has a non-zero ergotropy.
Figure 11: Power ($P_{\Delta W_{max}}$) of the total system for $0\ \leq\ \theta\ \leq\ \pi$ (x-axis) and $0\ \leq$ Iterations $\ \leq\ 30$ (y-axis); First row: Two-qubit systems where columns from left to right are case 1, 2, and 3 respectively. Brighter shades correspond to a greater value of power. Parameters: $\epsilon_{1j}=0\;\epsilon_{2B},\ \epsilon_{2j}=1\;\epsilon_{2B}$ for $j=A,B$, and $C$. $k_{B}T_{A}=4\;\epsilon_{2B}$ and $k_{B}T_{C}=0.4\;\epsilon_{2B}$. $\phi=0$. Second row: Same parameters as the first row but for three-qubit systems.
Figure 12: Same as figure 11 but for $\phi=\pi$.
Figures 11 and 12 follow a somewhat similar trend to that of figures 4 and  5 respectively (where Figs. 4 and  5 show $\Delta W_{max}$). Indeed, for the three qubit systems the maximum values of $P_{\Delta W_{max}}$ are reached when $\phi=\pi$ (Fig. 5) and areas with $\theta\approx\pi/2$, where initial coherence is large. When $\phi=0$ (Fig. 5) the regions of high power are achieved when $\theta\approx 0\ or\ \pi$, for which we have lower initial coherence. These maxima appear for similar values as the maxima of $\Delta W_{max}$, but at earlier iterations, see figs 4 and 5. Looking at eq. 12 we see that as iterations increase then a reduction in power is expected which is what we see in figures 11 and 12. This is why we see the power maxima at lower iterations than that of the maxima in figures 4 and  5. Directly comparing maxima for figures 11 and 12 we see a larger maximum power reached when $\phi=\pi$ with a value of $1.42\epsilon_{2B}$ compared with $1.07\epsilon_{2B}$ giving a $32.7\%$ increase.
@@
For completeness, in this appendix we present results for parameter scans with qubit B initialized with $\Phi=0$ and for the following quantities: Ergotropy of the total system (Fig. 13), Ergotropy ratio $W_{ratio}$ (Fig. 14), and Figure of merit FoM.(Fig. 15).
Figure 13: Ergotropy of the total system; Parameters: $\epsilon_{1j}=0\;\epsilon_{2B},\ \epsilon_{2j}=1\;\epsilon_{2B}$ for $j=A,B$, and $C$. $k_{B}T_{A}=4\;\epsilon_{2B}$ and $k_{B}T_{C}=0.4\;\epsilon_{2B}$. $\phi=0$. First row: Ergotropy of the total system for $0\ \leq\ \theta\ \leq\ \pi$ (x-axis) and $0\ \leq$ Iterations $\ \leq\ 30$ (y-axis) for two-qubit systems where columns from left to right are case 1, 2, and 3 respectively. Brighter shades correspond to a greater value of ergotropy. Second row: Same parameters as the first row but for three-qubit systems.
Figure 14: Ergotropy ratio $W_{ratio}$ for $\Phi=0$. Same parameters as figure 13.
Figure 15: Figure of merit FoM for $\Phi=0$. Same parameters as figure 13

### v2

@@ VIII Conclusions
EJF acknowledges support from EPSRC, grant number is EP/T518025/1; FSK acknowledges funding by the DFG within FOR 2724
@@
Figure 10 shows how changing $N$ affects the ergotropy for case 3 and the three-qubit system. We observe a “smoothing” of the evolution of the ergotropy for increasing $N$.
Figure 10: Ergotropy $W_{max}$ of the total system for $0\ \leq$ iterations $\ \leq\ 2N$. For each panel, $N$ is indicated on the x-axis; protocol corresponding to case 3 with three-qubit systems. Note that each panel corresponds to two cycles of iterations. Parameters: $\epsilon_{1j}=0,\ \epsilon_{2j}=1$ for $j=A,B$, and $C$.
@@
where $T_{\alpha}=\alpha/M$ is the process duration up to the $\alpha$ iteration and $M$ is the total number of iterations considered (here, $M=30$ corresponding to two cycles). For $\alpha=M$, $T_{\alpha}=1$. $\langle W_{\alpha}\rangle$ is the average work defined as $\langle W_{\alpha}\rangle=E^{\alpha}-E^{i}$, where $E^{\alpha}=Tr\left[\rho^{\alpha}H\right]$.
When evaluating the power of a quantum battery, there is more depth than just the difference in the average internal energy of the system. While the average internal energy plays a role, this is not a one-to-one comparison with the ergotropy, which as thoroughly discussed in the main article, being a better measure for quantum battery performance. With that in mind, we introduce,
@@
where $P_{\Delta W_{max}}$ is the power of charging or discharging and $\Delta W_{max}$ is the variation in ergotropy up to the $\alpha$ iteration. Equation (13) characterises the rate at which the system generates ergotropy. The total, ergotropy is not used as we are interested in charging from some initial state ($\rho^{i}$) that has a non-zero ergotropy.
Figure 11: Power ($P_{\Delta W_{max}}$) of the total system for $0\ \leq\ \theta\ \leq\ \pi$ (x-axis) and $0\ \leq$ iterations $\ \leq\ 30$ (y-axis); First row: Two-qubit systems where columns from left to right are cases 1, 2, and 3, respectively. Brighter shades correspond to a greater value of power. Parameters: $\epsilon_{1j}=0\epsilon_{2B},\ \epsilon_{2j}=1\epsilon_{2B}$ for $j=A,B$, and $C$. $k_{B}T_{A}=4\epsilon_{2B}$ and $k_{B}T_{C}=0.4\epsilon_{2B}$. $\phi=0$. Second row: The same parameters as the first row but for three-qubit systems.
Figure 12: The same as Figure 11, but for $\phi=\pi$.
[…]
@@
For completeness, in this Appendix, we present results for parameter scans with qubit B initialised with $\Phi=0$ and for the following quantities: The ergotropy of the total system (Figure 13), the ergotropy ratio $W_{ratio}$ (Figure 14), and the figure of merit FoM (Figure 15).
Figure 13: Ergotropy of the total system; parameters: $\epsilon_{1j}=0\epsilon_{2B},\ \epsilon_{2j}=1\epsilon_{2B}$ for $j=A,B$, and $C$. $k_{B}T_{A}=4\epsilon_{2B}$ and $k_{B}T_{C}=0.4\epsilon_{2B}$. $\phi=0$. First row: Ergotropy of the total system for $0\ \leq\ \theta\ \leq\ \pi$ (x-axis) and $0\ \leq$ iterations $\ \leq\ 30$ (y-axis) for two-qubit systems where columns from left to right are cases 1, 2, and 3, respectively. Brighter shades correspond to a greater value of ergotropy. Second row: The same parameters as the first row but for three-qubit systems.
Figure 14: Ergotropy ratio $W_{ratio}$ for $\Phi=0$, with the same parameters as Figure 13.
Figure 15: Figure of merit (FoM) for $\Phi=0$, with the same parameters as Figure 13.